# Benchmarking and Improving Large Vision-Language Models for Fundamental Visual Graph Understanding and Reasoning

### v2

@@ Abstract
Yingjie Zhu1,2, Xuefeng Bai1, Kehai Chen1, Yang Xiang2, Jun Yu1, Min Zhang1 1Harbin Institute of Technology, Shenzhen, China 2Peng Cheng Laboratory, Shenzhen, China zhuyj@stu.hit.edu.cn, {chenkehai,baixuefeng,zhangmin2021}@hit.edu.cn xiangy@pcl.ac.cn, yujun@hdu.edu.cn
@@ 1 Introduction
Graphs serve as a fundamental data structure across a wide range of domains, including social network analysis Schweimer et al. (2022), recommendation systems Zhang et al. (2023), knowledge graphs Zhang et al. (2024b), chemistry Cao et al. (2024) and semantic reasoning (Bai et al., 2022). Existing methods have achieved great success in enhancing understanding and reasoning abilities in graph-based tasks Kim et al. (2023a); Chen et al. (2024b). However, these approaches typically focus on specific graph types or tasks, posing challenges in designing versatile systems that are suitable for various tasks and graphs across diverse domains.
@@
Recently, Large Vision-Language Models (LVLMs) have exhibited outstanding performance across a wide range of downstream tasks by unifying various inputs in the form of images and processing them with human-like understanding and reasoning abilities Zhu et al. (2024); Zheng et al. (2024); Chen et al. (2024a). This triggers a growing interest in employing LVLMs for graph learning problems, as the vision modality offers a natural and intuitive way for comprehending structural information and facilitating general graph reasoning Poklukar et al. (2022). However, recent studies Wei et al. (2024); Li et al. (2024c); Ai et al. (2024) reveal significant challenges for LVLMs in graph-based learning tasks, where LVLMs achieve less than 15% accuracy on mathematical graph reasoning tasks, markedly below their performance in image and text reasoning Li et al. (2024c). Therefore, a natural challenge arises: why do LVLMs fail in graph learning, and how to enhance LVLMs to process graphs like professionals?
@@ 2.1 Graph Structure Generation
2.2 Tasks Design
  • Degree Query (DQ): Calculate the degree of the specified node.
@@ 3.5 Why LVLMs Fail on Fundamental Tasks?
In addition, by analyzing the generation results of LVLMs, we find that their main errors can be grouped into the following six categories: Relation Misunderstanding, Complexity Limitation, Structural Hallucination, Question-based Answering, Off-target Generation, and Format Error. Due to space constraint, we provide the detailed description in Appendix C.2.
@@ 4 The MCDGraph Framework
Table 5: Model performance (Acc/F1/EM_F1 for QA and F1 for FC) on various tasks. $\uparrow$ indicates an improvement compared to the original model. The complete experimental results are shown in Tab.17 and 17.
@@ 5.3 Results on Downstream Reasoning Tasks
FactKG We then evaluate MCDGraph on FactKG Kim et al. (2023b), a knowledge graph-based fact verification dataset collected from real-world data. As Tab.6 shows, MCDGraph gives consistently better performance than LVLMs on FactKG, suggesting that the proposed method can improve LVLMs in real-world graph-related tasks. Note that FactKG not only requires fundamental visual graph understanding and reasoning abilities but also relies on LVLMs’ understanding of semantic and logical relationships between entities, which our method does not address. Therefore, the improvement of LVLMs is not as significant as that of the graph theory problems. We also evaluate our method on two normal VQA tasks and the results are available in Appendix F.
@@ B.2 Evaluation Metrics
For the QA samples of NQ, NN, CR, RA, SRN, and NR tasks, we use (macro-averaged) F1 score and Hits@1 as in the previous QA benchmarks Rajpurkar et al. (2016); Zhang et al. (2018). For the QA samples of CP task, we employ EM_F1, which is the macro F1 score calculated based on the exact match between the predicted path and the ground truth path, and Label_Acc, which measures the accuracy of the model’s prediction on whether a path exists or not. For the QA samples of NNu, EN and DQ tasks, we compute the accuracy between predicted answers and ground truth. For the FC samples of all tasks, following Zhu et al. (2023) and Si et al. (2024) , we use macro F1 and accuracy as the metrics.
@@
Question-based Answering
@@ C.1 Comparison with Backbone LLMs
C.2 Error Analysis
We randomly select a batch of results generated by Qwen2-VL and InternVL2 on QA samples in VGCure and attribute their errors to the following six types:
  • Relation Misunderstanding: Failure to properly understand the relations between entities, leading to erroneous reasoning.
  • Complexity Limitation: When faced with intricate or highly complex graph structures, LVLMs struggle to process and understand the information effectively, often resulting in incomplete or erroneous outputs.
  • Structural Hallucination: Generating or perceiving structures that do not exist, leading to erroneous or misleading results that do not match the actual visual graph.
  • Question-based Answering: Directly using the entities mentioned in the question as answers, thus bypassing the need for deeper relation understanding or logical reasoning in the visual graph.
  • Off-target Generation: Deviation from the original task or misunderstanding of the question, leading to the generation of irrelevant answers.
  • Format Error: The output generated by the model is incorrectly formatted or unexpected.
Examples of each error are shown in the Tab.9.
@@ D.2 Implementation Details
Table 11: Model performance (Acc) on normal VQA tasks.
@@ E.1 Impact on Visual Styles
As the results shown in Tabs.12 and 13, although LVLMs never encounter the different style of visual graph during fine-tuning, our method can still improve the performance of the LVLMs on almost all tasks. This demonstrates the ability of our method to enhance LVLMs’ ability in capturing the structural information in visual graphs with excellent generalization. In addition, compared to Tab.5, it can be noticed that the experimental results before and after the change of style are similar, which indicates that the style itself has no effect on the evaluation of LVLMs.
@@ E.2 Impact of Naming Conventions
The results are shown in Tabs.14 and 15. We can observe that after altering the naming conventions, LVLMs continue to show similar trends on most of tasks. Therefore, the experimental analysis in the main text still holds. Meanwhile, the performance of LVLMs decreases on most of the tasks when confronted with different names. This might be due to the fact that the new naming conventions gives longer names to nodes and edges and recognizing these information increases the difficulty of the task. It is worth noting that in the face of the new naming conventions, our MCDGraph still improves the performance of LVLMs on most tasks, although the fine-tuning still uses the original naming conventions. This strongly demonstrates the effectiveness as well as the generalization of our method.
Appendix F Results on Normal VQA Tasks
We evaluate our MCDGraph on two normal VQA tasks, i.e., ScienceQA Lu et al. (2022) and AOKVQA Schwenk et al. (2022). According to the results shown in Tab.11, we can observe that after MCDGraph, LVLMs perform worse on these two VQA tasks. This is due to the fact that our MCDGraph is proposed for visual graph understanding and reasoning tasks, which are very different from normal VQA tasks. And it is worth noting that our method does not overly compromise the LVLMs’ performance on normal VQA tasks, which we consider acceptable
@@ Appendix G Case Study
Fig.20 illustrates a case on Nested Relation Query task. We can observe that Qwen2-VL, LLaVA-13B and mPLUG-Owl3 all make the error of Question-based Answering when confronted with this question, i.e., using “R1”, “E9”, and “E3” mentioned in the question as the generated answer. InternLM-XC2.5 makes the error of Relation Misunderstanding, i.e., the edge pointing from E3 to E1 is also used as the answer. However, Qwen2-VL after applying MCDGraph can answer this question correctly. This demonstrates that the proposed method can improve the fundamental graph structure understanding of LVLMs, thus avoiding the occurrence of the previously mentioned errors.

### v3

@@ Abstract
Yingjie Zhu1,2, Xuefeng Bai1111Corresponding Author, Kehai Chen1,2, Yang Xiang2111Corresponding Author, Jun Yu3, Min Zhang1,2 1Institute of Computing and Intelligence, Harbin Institute of Technology, Shenzhen, China 2Peng Cheng Laboratory, Shenzhen, China 3School of Intelligence Science and Engineering, Harbin Institute of Technology, Shenzhen, China {baixuefeng,chenkehai,yujun,zhangmin2021}@hit.edu.cn zhuyj@stu.hit.edu.cn, xiangy@pcl.ac.cn
@@ 1 Introduction
Graphs serve as a fundamental data structure across a wide range of domains, including social network analysis Schweimer et al. (2022), recommendation systems Zhang et al. (2023), knowledge graphs Zhang et al. (2024b), chemistry Cao et al. (2024), biomedical molecules Liu et al. (2023) and semantic reasoning (Bai et al., 2022). Existing methods have achieved great success in enhancing understanding and reasoning abilities in graph-based tasks Kim et al. (2023a); Chen et al. (2024). However, these approaches typically focus on specific graph types or tasks, posing challenges in designing versatile systems that are suitable for various tasks and graphs across diverse domains.
@@
Recently, Large Vision-Language Models (LVLMs) have exhibited outstanding performance across a wide range of downstream tasks by unifying various inputs in the form of images and processing them with human-like understanding and reasoning abilities Zhu et al. (2024); Zheng et al. (2024); Zhang et al. (2025). This triggers a growing interest in employing LVLMs for graph learning problems, as the vision modality offers a natural and intuitive way for comprehending structural information and facilitating general graph reasoning Poklukar et al. (2022). However, recent studies Wei et al. (2024); Li et al. (2024c); Ai et al. (2024) reveal significant challenges for LVLMs in graph-based learning tasks, where LVLMs achieve less than 15% accuracy on mathematical graph reasoning tasks, markedly below their performance in image and text reasoning Li et al. (2024c). Therefore, a natural challenge arises: why do LVLMs fail in graph learning, and how to enhance LVLMs to process graphs like professionals?
@@ 2.1 Graph Structure Generation
2.2 Tasks Design
  • Degree Query (DQ): Calculate the degree of the specified node.
@@ 3.5 Why LVLMs Fail on Fundamental Tasks?
Error Type R.M. C.L. S.H. E.A. Ot.G. F.E.
Percentage 54% 7% 17% 9% 5% 8%
Table 5: Frequency of each error type.
3.6 Error Analysis
In addition, by analyzing a batch of results generated by Qwen2-VL and InternVL2 on QA samples in VGCure, we find that their main errors can be grouped into the following six categories:
  • Relation Misunderstanding (R.M.): Failure to properly understand the relations between entities, leading to erroneous reasoning.
  • Complexity Limitation (C.L.): When faced with intricate or highly complex graph structures, LVLMs struggle to process and understand the information effectively, often resulting in incomplete or erroneous outputs.
  • Structural Hallucination (S.H.): Generating or perceiving structures that do not exist, leading to erroneous or misleading results that do not match the actual visual graph.
  • Entity-based Answering (E.A): Directly using the entities mentioned in the question as answers, thus ignoring deeper relation understanding or logical reasoning in the visual graph.
  • Off-target Generation (Ot.G.): Deviation from the task or misunderstanding of the question, leading to the generation of irrelevant answers.
  • Format Error (F.E.): The output of the model is incorrectly formatted or unexpected.
Examples of each error are shown in the Tab.11.
Error Distribution Analysis We also calculate the overall distribution of each error. The results are shown in Tab.5. It can be observed that, (I) Relation Misunderstanding appears most frequently due to LVLMs’ limited capacity to effectively capture structural and relational information in the visual graph. (II) Structural Hallucination also appears frequently due to the inherent hallucination tendency of LLMs. (III) Despite demonstrating robust instruction-following capabilities, LVLMs remain prone to errors like Off-target Generation and Format Error. (IV) Complexity Limitation and Entity-based Answering also account for a certain percentage, i.e., LVLMs may be stumped when faced with overly complex visual graphs, or directly use the entities mentioned in the question as answers.
Primary Error Pattern Analysis We further analyze the primary error pattern of each task. We find that, (I) Generally, the primary error occurring in all tasks except the Node Number Query is Relation Misunderstanding. (II) Structural Hallucination occurs more frequently in Nested Node Query, Conjunctive Relation Query, Relation Analogy Query, Shared Relation Neighbor Query, and Nested Relation Query tasks. (III) Format Error is usually found in Connective Path Query task.
@@ 4 The MCDGraph Framework
Table 6: Model performance (Acc/F1/EM_F1 for QA and F1 for FC) on various tasks. $\uparrow$ indicates an improvement compared to the original model. The complete experimental results are shown in Tab.19 and 19.
@@ 5.2 Results and Analysis
Ablation Study We also conduct ablation study on QA samples with Qwen2-VL and the results are shown in Tab.7. It can be observed that after removing different training tasks, although the LVLM may have a slight performance improvement on some tasks, it may lose certain capabilities causing its performance to plummet on some task (as evidenced by the red results in the Tab.7). Meanwhile, all three self-supervised tasks enhance the LVLMs’ ability to capture structural and relational information from different dimensions, complementing each other in order to comprehensively improve the LVLM’s overall performance on all tasks.
Models Understanding Reasoning
NNu EN DQ NQ NN CR RA SRN NR CP
Qwen2-VL 97.80 16.38 48.09 16.18 16.52 21.02 14.19 19.48 12.73 12.90
w MCDGraph 98.34 25.92 60.94 25.44 13.32 26.14 13.14 20.74 14.44 11.95
- Masked 99.02 32.11 57.75 22.84 14.15 23.29 15.32 19.15 5.01 24.61
- Contrastive 99.64 28.70 64.02 25.35 15.16 25.13 11.83 21.11 15.20 5.33
- Description 2.00 13.27 57.80 23.09 14.14 23.47 10.16 19.87 16.99 8.41
Table 7: Ablation study on across various tasks.
@@ 5.3 Results on Downstream Reasoning Tasks
FactKG We then evaluate MCDGraph on FactKG Kim et al. (2023b), a knowledge graph-based fact verification dataset collected from real-world data. As Tab.8 shows, MCDGraph gives consistently better performance than LVLMs on FactKG, suggesting that the proposed method can improve LVLMs in real-world graph-related tasks. Note that FactKG not only requires fundamental visual graph understanding and reasoning abilities but also relies on LVLMs’ understanding of semantic and logical relationships between entities, which our method does not address. Therefore, the improvement of LVLMs is not as significant as that of the graph theory problems. We also evaluate our method on two general vision reasoning tasks and the results are available in Appendix F.
@@ 6 Related Work
Graph Benchmarks for GNNs Rozemberczki et al. (2021) construct Wikipedia-based graphs with pages as nodes and hyperlinks as edges. Hu et al. (2020) present realistic datasets spanning social, biological, molecular, code, and knowledge graphs. Mernyei and Cangea (2020) propose Wiki-CS, a GNN benchmark based on Computer Science articles. Morris et al. (2020) release TUDataset, covering 120 datasets across multiple domains for graph classification and regression. To address homophily limitations, Lim et al. (2021) introduce large-scale, non-homophilous graph datasets. Dwivedi et al. (2022) propose the Long Range Graph Benchmark to evaluate models on long-range interaction reasoning tasks. Different from these graph benchmarking approach, the proposed VGCure focuses on exploring the fundamental graph understanding and reasoning capabilities of LVLMs, showcasing their potential to unify multimodal information processing through a unified visual learning paradigm.
@@ 7 Conclusion
Acknowledgments
We would like to thank the anonymous reviewers and meta-reviewer for their insightful suggestions. This work was partly supported by the National Natural Science Foundation of China (62406091, 62276077, 62125201, U23B2055, U24A20328, U24B20174, 62350710797), the National Science and Technology Innovation 2030 Major program (2024ZD01NL00101), and Shenzhen Science and Technology Program (KQTD2024072910215406, ZDSYS20230626091203008).
@@ B.2 Evaluation Metrics
For the QA samples of NQ, NN, CR, RA, SRN, and NR tasks, we use (macro-averaged) F1 score and Hits@1 as in the previous QA benchmarks Rajpurkar et al. (2016); Zhang et al. (2018). For the QA samples of CP task, we employ EM_F1, which is the macro F1 score calculated based on the exact match between the predicted path and the ground truth path, and Label_Acc, which measures the accuracy of the model’s prediction on whether a path exists or not. For the QA samples of NNu, EN and DQ tasks, we compute the accuracy between predicted answers and ground truth. For the FC samples of all tasks, following Si et al. (2024a, b), we use macro F1 and accuracy as the metrics.
@@
Entity-based Answering
@@ D.2 Implementation Details
Table 13: Model performance (Acc) on general VQA tasks.
@@ E.1 Impact on Visual Styles
As the results shown in Tabs.14 and 15, although LVLMs never encounter the different style of visual graph during fine-tuning, our method can still improve the performance of the LVLMs on almost all tasks. This demonstrates the ability of our method to enhance LVLMs’ ability in capturing the structural information in visual graphs with excellent generalization. In addition, compared to Tab.6, it can be noticed that the experimental results before and after the change of style are similar, which indicates that the style itself has no effect on the evaluation of LVLMs.
@@ E.2 Impact of Naming Conventions
The results are shown in Tabs.16 and 17. We can observe that after altering the naming conventions, LVLMs continue to show similar trends on most of tasks. Therefore, the experimental analysis in the main text still holds. Meanwhile, the performance of LVLMs decreases on most of the tasks when confronted with different names. This might be due to the fact that the new naming conventions gives longer names to nodes and edges and recognizing these information increases the difficulty of the task. It is worth noting that in the face of the new naming conventions, our MCDGraph still improves the performance of LVLMs on most tasks, although the fine-tuning still uses the original naming conventions. This strongly demonstrates the effectiveness as well as the generalization of our method.
Appendix F Results on General VQA Tasks
We evaluate our MCDGraph on two general VQA tasks, i.e., ScienceQA Lu et al. (2022) and AOKVQA Schwenk et al. (2022). According to the results shown in Tab.13, we can observe that after MCDGraph, LVLMs perform worse on these two VQA tasks. This is due to the fact that our MCDGraph is proposed for visual graph understanding and reasoning tasks, which are very different from general VQA tasks. And it is worth noting that our method does not overly compromise the LVLMs’ performance on general VQA tasks, which we consider acceptable.
@@ Appendix G Case Study
Fig.20 illustrates a case on Nested Relation Query task. We can observe that Qwen2-VL, LLaVA-13B and mPLUG-Owl3 all make the error of Entity-based Answering when confronted with this question, i.e., using “R1”, “E9”, and “E3” mentioned in the question as the generated answer. InternLM-XC2.5 makes the error of Relation Misunderstanding, i.e., the edge pointing from E3 to E1 is also used as the answer. However, Qwen2-VL after applying MCDGraph can answer this question correctly. This demonstrates that the proposed method can improve the fundamental graph structure understanding of LVLMs, thus avoiding the occurrence of the previously mentioned errors.